You are trying to save money, so you decide to stick $5 a day under your pillow. At the end of the year, you deposit the lump sum. How much will you have after 20 years if the assumed rate is 5%?
Answer #1

deposit end of each year = number of days in a year*amount stuck under pillow each day

=365*5=1825

amount in 20 years

FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
FV= 1825*(((1+ 5/100)^20-1)/(5/100))
FV = 60345.37